OpenStudy (anonymous):

Suppose you bought a 100-foot fence that you wanted to put around a square section of yard that will be used as a garden. The side length of your garden is 13 feet. The fencing was on sale for $12.60 per foot.How much fencing is left over?

OpenStudy (anonymous):

math o.0 lol

OpenStudy (anonymous):

give me a second lol ill help :)

OpenStudy (anonymous):

sorry im new at this!!

OpenStudy (anonymous):

its cool..i did the same thin >.< im no good at math and the person i was gonna tag isnt online :( im sorry

OpenStudy (anonymous):

its okayy lol

OpenStudy (anonymous):

@redred413 @dmezzullo @zaphodplaysitsafe

OpenStudy (anonymous):

lol mybe theyll help and tank yew for zee madal

OpenStudy (anonymous):

you welcomee

OpenStudy (anonymous):

fencing left = total length of fence - total perimeter of square garden = 100 - (4*13) = 100 - 52 = 48 feet

OpenStudy (anonymous):

thankss :)